OpenStudy (anonymous):

determine the coordinates of the points on the curve y=tanx between 0 and 2pi where the tangent line is perpendcular to y=-1/2x+3

OpenStudy (anonymous):

@undeadknight26 @Callisto

OpenStudy (anonymous):

tanx become secx^2 right?

undeadknight26 (undeadknight26):

*head be spinning* Umm...Umm Blue? @aaronq @sleepyjess @dumbcow @Data_LG2 @dan815

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (irishboy123):

yes, y ' = sec^2 x

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (dumbcow):

slope perpendicular to -1/2 is 2 set derivative equal to 2 \[\sec^2 x = 2\] \[\sec x = \pm \sqrt{2}\] \[\cos x = \pm \frac{1}{\sqrt{2}}\] \[x = \frac{\pi}{4} + \frac{\pi}{2}k\]